Weight of 50 atoms of an element is 3000 amu, calculate no. Of atoms in 240gm of the elements. 
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
 
Mass of 1 atom = 3000/ 50 = 60 u
 
mass of 1 mol of atoms = 60 gm
 
number of moles = 240/60 = 4 mol
 
hence number of atoms = 4 * 6.023 * 1023 = 24.092 * 1023 atoms